### v3

# Derivative-Free Guidance in Continuous and Discrete Diffusion Models with Soft Value-Based Decoding

Abstract
Diffusion models excel at capturing the natural design spaces of images, molecules, DNA, RNA, and protein sequences. However, rather than merely -generating designs that are natural, we often aim to optimize downstream reward functions while preserving the naturalness of these design spaces. Existing methods for achieving this goal often require “differentiable” proxy models (e.g., classifier guidance or DPS) or involve computationally expensive fine-tuning of diffusion models (e.g., classifier-free guidance, RL-based fine-tuning). In our work, we propose a new method to address these challenges. Our algorithm is an iterative sampling method that integrates soft value functions, which looks ahead to how intermediate noisy states lead to high rewards in the future, into the standard inference procedure of pre-trained diffusion models. Notably, our approach avoids fine-tuning generative models and eliminates the need to construct differentiable models. This enables us to (1) directly utilize non-differentiable features/reward feedback, commonly used in many scientific domains, and (2) apply our method to recent discrete diffusion models in a principled way. Finally, we demonstrate the effectiveness of our algorithm across several domains, including image generation, molecule generation, and DNA/RNA sequence generation. The code is available at https://github.com/masa-ue/SVDD.
1 Introduction
Diffusion models have gained popularity as powerful generative models. Their applications extend beyond image generation to include natural language generation (Sahoo et al., 2024; Shi et al., 2024; Lou et al., 2023), molecule generation (Jo et al., 2022; Vignac et al., 2022), and biological (DNA, RNA, protein) sequence generation (Avdeyev et al., 2023; Stark et al., 2024), and in each of these domains diffusion models have been shown to be very effective at capturing complex natural distributions. However, in practice, we might not only want to generate realistic samples, but to produce samples that optimize specific downstream reward functions while preserving naturalness by leveraging pre-trained models. For example, in computer vision, we might aim to generate natural images with high aesthetic and alignment scores. In drug discovery, we may seek to generate valid molecules with high QED/SA/docking scores (Lee et al., 2023; Jin et al., 2018) or natural RNAs (such as mRNA vaccines (Cheng et al., 2023)) with high translational efficiency and stability (Castillo-Hair and Seelig, 2021; Asrani et al., 2018), and natural DNAs with high cell-specificity (Gosai et al., 2023; Taskiran et al., 2024; Lal et al., 2024).
The optimization of downstream reward functions using pre-trained diffusion models has been approached in various ways. In our work, we focus on non-fine-tuning-based methods because fine-tuning generative models (e.g., when using classifier-free guidance (Ho et al., 2020) or RL-based fine-tuning (Black et al., 2023; Fan et al., 2023; Uehara et al., 2024; Clark et al., 2023; Prabhudesai et al., 2023)) often becomes computationally intensive, especially as pre-trained generative models grow larger in the era of “foundation models”. Although classifier guidance and its variants (e.g., Dhariwal and Nichol (2021); Song et al. (2020); Chung et al. (2022); Bansal et al. (2023); Ho et al. (2022)) have shown success as non-fine-tuning methods in these settings, they face significant challenges. Firstly, as they would require constructing differentiable proxy models, they cannot directly incorporate useful domain-specific non-differentiable features (e.g., molecular/protein descriptors (van Westen et al., 2013; Ghiringhelli et al., 2015; Gainza et al., 2020)) or non-differentiable reward feedback (e.g., physics-based simulations such as Vina and Rosetta (Trott and Olson, 2010; Alhossary et al., 2015; Alford et al., 2017)). This limitation also hinders the application of current classifier guidance methods to recently developed discrete diffusion models (Austin et al., 2021; Campbell et al., 2022; Sahoo et al., 2024; Shi et al., 2024; Lou et al., 2023) in a principled manner (i.e., without transforming the discrete space into the Euclidean space). Additionally, gradient computation can be memory-intensive with large proxy models.
Figure 1: Summary of SVDD. The notation $v$ denotes value functions that predict reward $r(x_{0})$ (at time 00) from states at time $t-1$. SVDD involves two steps: (1) generating multiple noisy states from pre-trained models and (2) selecting the state with the highest value according to the value function.
We propose a novel method, SVDD (Soft Value-based Decoding in Diffusion models), for optimizing downstream reward functions in diffusion models (summarized in Figure 1) to address the aforementioned challenges. Inspired by recent literature on RL-based fine-tuning (Uehara et al., 2024), we first introduce soft value functions that serve as look-ahead functions, indicating how intermediate noisy samples lead to high rewards in the future of the diffusion denoising process. After learning (or approximating) these value functions, we present a new inference-time technique, SVDD, which obtains multiple noisy states from the policy (i.e., denoising map) of pre-trained diffusion models and selects the sample with the highest value function at each time step. Specifically, we introduce two algorithms (SVDD-MC and SVDD-PM) depending on how we estimate value functions. Notably, the SVDD-PM approach, leveraging the characteristics of diffusion models (e.g., utilizing the forward process in diffusion models to directly map $t$ to 00 in terms of expectation in Figure 1), requires no additional training as long as we have access to the reward feedback.
Table 1: A comparison of our method with a number of prior approaches. “Non-differentiable” refers to the method’s capability to operate without requiring differentiable proxy models. “No Training” means that no additional training of the diffusion model is required as long as we have access to the reward feedback. Regarding the comparison with standard SMC-based methods, refer to Section A.
No fine-tuning Non-differentiable No Training
Classifier guidance (Dhariwal and Nichol, 2021) $\checkmark$
DPS (Chung et al., 2022) $\checkmark$ $\checkmark$
Classifier-free (Ho and Salimans, 2022) $\checkmark$
RL fine-tuning $\checkmark$
SVDD-MC (Ours) $\checkmark$ $\checkmark$
SVDD-PM (Ours) $\checkmark$ $\checkmark$ $\checkmark$
Our contributions are summarized as follows (also see Table 1). We propose a novel technique for optimizing downstream reward functions in pre-trained diffusion models that eliminates the need to construct differentiable proxy models and avoids the need to fine-tune the generative model itself. The first property allows for the use of non-differentiable reward feedback, which is common in many scientific fields, and makes our method applicable to recent discrete diffusion models without any modification. The second property addresses the high computational cost associated with fine-tuning diffusion models. We demonstrate the effectiveness of our methods across various domains, including image generation, molecule generation, and DNA/RNA generation.
2 Related Works
Here, we summarize related work. We first outline methods relevant to our goal, categorizing them based on whether they involve fine-tuning. We then discuss related directions, such as discrete diffusion models, where our method excels, and decoding in autoregressive models.
Non-fine-tuning methods.
There are primarily two methods for optimizing downstream functions in diffusion models without fine-tuning. Regarding the comparison with the most related works based on sequential Monte Carlo (SMC), we defer the discussion to Section 5.4 and Section A.
  • Classifier guidance (Dhariwal and Nichol, 2021; Song et al., 2020): It has been widely used to condition pre-trained diffusion models without fine-tuning. Although these methods do not originally focus on optimizing reward functions, they can be applied for this purpose (Uehara et al., 2024, Section 6.2). In this approach, an additional derivative of a certain value function is incorporated into the drift term (mean) of pre-trained diffusion models during inference. Subsequent variants (e.g., Chung et al. (2022); Ho et al. (2022); Bansal et al. (2023); Guo et al. (2024); Wang et al. (2022); Yu et al. (2023)) have been proposed to simplify the learning of value functions. However, these methods require constructing differentiable models, which limits their applicability to non-differentiable features/reward feedbacks commonly encountered in scientific domains as mentioned in Section 1. Additionally, this approach cannot be directly extended to discrete diffusion models in a principle way  [^1] We note that a notable exception has been recently proposed (Nisonoff et al., 2024). However, our method can be applied to both continuous and discrete diffusion models in a unified manner. Furthermore, their method in practice requires the differentiability of proxy models, unlike our proposal. Regarding more details, refer to Section B. . Our approach aims to address these challenges.
  • Best-of-N: The naive way is to generate multiple samples and select the top samples based on the reward functions, known as best-of-N in the literature on (autoregressive) LLMs (Stiennon et al., 2020; Nakano et al., 2021; Touvron et al., 2023; Beirami et al., 2024; Gao et al., 2023). This approach is significantly less efficient than ours, as our method uses soft-value functions that predict how intermediate noisy samples lead to high rewards in the future. We validate this experimentally in Section 6.
Fine-tuning of diffusion models.
Several methods exist for fine-tuning generative models to optimize downstream reward functions, such as classifier-free guidance (Ho and Salimans, 2022) and RL-based fine-tuning (Fan et al., 2023; Black et al., 2023)/its variants (Dong et al., 2023). However, these approaches often come with caveats, including high computational costs and the risk of easily forgetting pre-trained models. In our work, we propose an inference-time technique that eliminates the need for fine-tuning generative models, similar to classifier guidance.
Discrete diffusion models.
Based on seminal works Austin et al. (2021); Campbell et al. (2022), recent work on masked diffusion models (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024) has demonstrated their strong performance in natural language generation. Additionally, they have been applied to biological sequence generation (e.g., DNA, protein sequences in Campbell et al. (2024); Sarkar et al. (2024)). In these cases, the use of diffusion models over autoregressive models is particularly apt, given that many biological sequences ultimately adopt complex three-dimensional structures. We also note that ESM3 (Hayes et al., 2024), a widely recognized foundational model in protein sequence generation, bears similarities to masked diffusion models.
Despite its significance, it cannot be integrated with standard classifier guidance because adding a continuous gradient to a discrete objective is not inherently valid. Unlike standard classifier guidance, our algorithm can be seamlessly applied to discrete diffusion models.
Decoding in autoregressive models with rewards.
The decoding strategy, which dictates how sentences are generated from the model, is a critical component of text generation in autoregressive language models (Wu et al., 2016; Chorowski and Jaitly, 2016; Leblond et al., 2021). Recent studies have explored inference-time techniques for optimizing downstream reward functions Dathathri et al. (2019); Yang and Klein (2021); Qin et al. (2022); Mudgal et al. (2023); Zhao et al. (2024); Han et al. (2024). While there are similarities between these works and ours, to the best of our knowledge, no prior work has extended such methodologies to diffusion models. Furthermore, our approach leverages characteristics unique to diffusion models that are not present in autoregressive models such as SVDD-PM.
3 Preliminaries and Goal
In this section, we describe the standard method for training diffusion models and outline the objective of our work: optimizing downstream reward functions given pre-trained diffusion models.
3.1 Diffusion Models
In diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020), our goal is to learn a sampler $p(x)\in\Delta(\mathcal{X})$ given data consisting of $(x)\in\mathcal{X}$. The training process for a standard diffusion model is summarized as follows. First, we introduce a (fixed) forward process $q_{t}:\mathcal{X}\to\Delta(\mathcal{X})$. Now, after introducing the forward process, we aim to learn a backward process: $\{p_{t}\}$ where each $p_{t}$ is $\mathcal{X}\to\Delta(\mathcal{X})$ so that the induced distributions induced by the forward process and backward process match marginally. For this purpose, by parametrizing the backward processes with $\theta\in\mathbb{R}^{d}$, we typically use the following loss function:
$\mathbb{E}_{x_{1},\cdots,x_{T}\sim q(\cdot|x_{0})}\left[-\log p_{0}(x_{0}|x_{1})+\sum_{t=1}^{T}\mathrm{KL}(q_{t}(\cdot\mid x_{t-1})\|p_{t}(\cdot\mid x_{t+1};\theta))+\mathrm{KL}(q_{T}(\cdot)\|p_{T}(\cdot))\right],$
which is derived from the variational lower bound of the negative likelihood (i.e., ELBO).
Here are two examples of concrete parameterizations. Let $\alpha_{t}\in\mathbb{R}$ be a noise schedule.
Example 1 (Continuous space).
When $\mathcal{X}$ is Euclidean space, we typically use the Gaussian distribution $q_{t}(\cdot\mid x)=\mathcal{N}(\sqrt{\alpha_{t}}x,(1-\alpha_{t}))$. Then, the backward process is parameterized as
$\mathcal{N}\left(\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t+1})x_{t}+\sqrt{\alpha_{t-1}}(1-\alpha_{t})\hat{x}_{0}(x_{t};\theta)}{1-\bar{\alpha}_{t}},\frac{(1-\alpha_{t})(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\right),$
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. Here, $\hat{x}_{0}(x_{t};\theta)$ is a neural network that predicts $x_{0}$ from $x_{t}$ (i.e., $\mathbb{E}_{q}[x_{0}|x_{t}]$).
Example 2 (Discrete space in Sahoo et al. (2024); Shi et al. (2024)).
Let $\mathcal{X}$ be a space of one-hot column vectors $\{x\in\{0,1\}^{K}:\sum_{i=1}^{K}x_{i}=1\}$, and $\mathrm{Cat}(\pi)$ be the categorical distribution over $K$ classes with probabilities given by $\pi\in\Delta^{K}$ where $\Delta^{K}$ denotes the K-simplex. A typical choice is $q_{t}(\cdot\mid x)=\mathrm{Cat}(\alpha_{t}x+(1-\alpha_{t})\mathbf{m})$ where $\mathbf{m}=[0,\cdots,0,\mathrm{Mask}]$. Then, the backward process is parameterized as
$\begin{cases}\mathrm{Cat}(x_{t}),\quad\mathrm{if}\,x_{t}\neq\mathbf{m}\\ \mathrm{Cat}\left(\frac{(1-\alpha_{t-1})\mathbf{m}+(\alpha_{t-1}-\alpha_{t})\hat{x}_{0}(x_{t};\theta)}{1-\alpha_{t}}\right),\quad\mathrm{if}\,x_{t}=\mathbf{m}.\end{cases}$
Here, $\hat{x}_{0}(x_{t};\theta)$ is a neural network that predicts $x_{0}$ from $x_{t}$. Note that when considering a sequence of $L$ tokens ($x^{1:L})$, we use the direct product: $p_{t}(x^{1:L}_{t}|x^{1:L}_{t+1})=\prod_{l=1}^{L}p_{t}(x^{l}_{t}|x^{1:L}_{t+1})$.
After learning the backward process, we can sample from a distribution that emulates training data distribution (i.e., $p(x)$) by sequentially sampling $\{p_{t}\}_{t=T}^{0}$ from $t=T$ to $t=0$.
Notation.
The notation $\delta_{a}$ denotes a Dirac delta distribution centered at $a$. The notation $\propto$ indicates that the distribution is equal up to a normalizing constant. With slight abuse of notation, we often denote $p_{T}(\cdot|\cdot,\cdot)$ by $p_{T}(\cdot)$.
3.2 Objective: Generating Samples with High Rewards While Preserving Naturalness
We consider a scenario where we have a pre-trained diffusion model, which is trained using the loss function explained in Section 3.1. These pre-trained models are typically designed to excel at characterizing the natural design space (e.g., image space, biological space, or chemical space) by emulating the extensive training dataset. Our work focuses on obtaining samples that also optimize downstream reward functions $r:\mathcal{X}\to\mathbb{R}$ (e.g., QED and SA in molecule generation) while maintaining the naturalness by leveraging pre-trained diffusion models. We formalize this goal as follows.
[…]
$p^{(\alpha)}(x)$ $:=\mathop{\mathrm{argmax}}_{p\in[\Delta(\mathcal{X})]}\underbrace{\mathbb{E}_{x\sim p(\cdot)}[r(x)]}_{\textbf{term (a)}}-\alpha\underbrace{\mathrm{KL}(p(\cdot)\|p^{\mathrm{pre}}(\cdot))}_{\textbf{term(b)}}$
$\propto\exp(r(x)/\alpha)p^{\mathrm{pre}}(x).$
Here, the term (a) is introduced to optimize reward function, while the term (b) is used to maintain the naturalness of the generated samples.
Existing methods.
Several existing approaches target this goal (or its variant), as discussed in Section 2, including classifier guidance, fine-tuning (RL-based or classifier-free), and Best-of-N. In our work, we focus on non-fine-tuning-based methods, such as classifier guidance and Best-of-N. Specifically, we aim to address the limitations of these methods: the requirement for differentiable proxy models in classifier guidance and the inefficiency of Best-of-N.
Remark 1.
We can easily extend all the diffusion models discussed in our paper to cases where the pre-trained model is a conditional diffusion model.
4 Soft Value-Based Decoding in Diffusion Models
In this section, we first present the motivation behind developing our new algorithm. We then introduce our algorithm, which satisfies desired properties, i.e., the lack of need for fine-tuning or constructing differentiable models.
4.1 Key Observation
We introduce several key concepts. First, for $t\in[T+1,\cdots,1]$, we define the soft value function:
$v_{t-1}(\cdot):=\alpha\log\mathbb{E}_{x_{0}\sim p^{\mathrm{pre}}(x_{0}|x_{t-1})}\left[\exp\left(\frac{r(x_{0})}{\alpha}\right)|x_{t-1}=\cdot\right],$
where $\mathbb{E}_{\{p^{\mathrm{pre}}\}}[\cdot]$ is induced by $\{p^{\mathrm{pre}}_{t}(\cdot|x_{t+1})\}_{t=T}^{0}$. This value function represents the expected future reward at $t=0$ from the intermediate noisy state at $t-1$. Next, we define the following soft optimal policy (denoising process) $p^{\star,\alpha}_{t-1}:\mathcal{X}\to\Delta(\mathcal{X})$ weighted by value functions $v_{t-1}:\mathcal{X}\to\mathbb{R}$:
[…]
Here, we refer to $v_{t}$ as soft value functions, $p^{\star,\alpha}_{t}$ as soft optimal policies, respectively, because they literally correspond to soft value functions and soft optimal policies where we embed diffusion models into entropy-regularized MDPs, as demonstrated in Uehara et al. (2024).
With this preparation in mind, we utilize the following key observation:
Theorem 1 (From Theorem 1 in Uehara et al. (2024)).
The distribution induced by $\{p^{\star,\alpha}_{t}(\cdot|x_{t+1})\}_{t=T}^{0}$ is the target distribution $p^{(\alpha)}(x)$, i.e.,
[…]
While Uehara et al. (2024) presents this theorem, they use it primarily to interpret their RL-based fine-tuning introduced in Fan et al. (2023); Black et al. (2023). In contrast, our work explores how to convert this into a new fine-tuning-free optimization algorithm.
Our motivation for a new algorithm.
Theorem 1 states that if we can hypothetically sample from $\{p^{\star,\alpha}_{t}\}_{t=T}^{0}$, we can sample from the target distribution $p^{(\alpha)}$. However, there are two challenges in sampling from each $p^{\star,\alpha}_{t-1}$:
  1. 1.
    The soft-value function $v_{t-1}$ in $p^{\star,\alpha}_{t-1}$ is unknown.
  2. 2.
    It is unnormalized (i.e., calculating the normalizing constant is often hard).
We address the first challenge later in Section 4.3. Assuming the first challenge is resolved, we consider how to tackle the second challenge. A natural approach is to use importance sampling (IS):
$p^{\star,\alpha}_{t-1}(\cdot|x_{t},c)\approx\sum_{m=1}^{M}\frac{w^{\langle m\rangle}_{t-1}}{\sum_{j}w^{\langle j\rangle}_{t-1}}\delta_{x^{\langle m\rangle}_{t-1}},\quad\{x^{\langle m\rangle}_{t-1}\}_{m=1}^{M}\sim p^{\mathrm{pre}}_{t-1}(\cdot\mid x_{t}),$
where $w^{\langle m\rangle}_{t-1}:=\exp(v_{t}(x^{\langle m\rangle}_{t-1})/\alpha)$. Thus, we can approximately sample from $p^{\star,\alpha}_{t-1}(\cdot|x_{t})$ by obtaining multiple ($M$) samples from pre-trained diffusion models and selecting the sample based on an index, which is determined by sampling from the categorical distribution with mean $\{w^{\langle m\rangle}_{t-1}/\sum_{j}w^{\langle j\rangle}_{t-1}\}_{m=1}^{\langle M\rangle}$.
Difference over Best-of-N.
Best-of-N, which generates multiple samples and selects the highest reward sample, is technically considered IS where the proposal distribution is the entire $p^{\mathrm{pre}}(x_{0})=\int\prod_{t}\{p^{\mathrm{pre}}_{t}(x_{t-1}\mid x_{t})\}dx_{1:T}$. However, the use of importance sampling in our algorithm differs significantly, as we apply it at each time step to approximate each soft-optimal policy.
4.2 Inference-Time Algorithm
Algorithm 1 SVDD (Soft Value-Based Decoding in Diffusion Models)
1:  Require: Estimated soft value function $\{\hat{v}_{t}\}_{t=T}^{0}$ (refer to Algorithm 2 or Algorithm 3), pre-trained diffusion models $\{p^{\mathrm{pre}}_{t}\}_{t=T}^{0}$, hyperparameter $\alpha\in\mathbb{R}$
2:  for $t\in[T+1,\cdots,1]$ do
3:     Get $M$ samples from pre-trained polices $\{x^{\langle m\rangle}_{t-1}\}_{m=1}^{M}\sim p^{\mathrm{pre}}_{t-1}(\cdot|x_{t})$, and for each $m$, calculate $w^{\langle m\rangle}_{t-1}:=\exp(\hat{v}_{t-1}(x^{\langle m\rangle}_{t-1})/\alpha)$
4:     $x_{t-1}\leftarrow x^{\langle\zeta_{t-1}\rangle}_{t-1}$ after selecting an index: $\zeta_{t-1}\sim\mathrm{Cat}\left(\left\{\frac{w^{\langle m\rangle}_{t-1}}{\sum_{j=1}^{M}w^{\langle j\rangle}_{t-1}}\right\}_{m=1}^{M}\right),\,$
5:  end for
6:  Output: $x_{0}$
Now, we introduce our algorithm, which leverages the observation, in Algorithm 1. Our algorithm is an iterative sampling method that integrates soft value functions into the standard inference procedure of pre-trained diffusion models. Each step is designed to approximately sample from a value-weighted policy $\{p^{\star,\alpha}_{t}\}_{t=T}^{0}$.
We have several important things to note.
  • When $\alpha=0$, Line 4 corresponds to $\zeta_{t-1}=\mathop{\mathrm{argmax}}_{m\in[1,\cdots,M]}\hat{v}_{t-1}(x^{\langle m\rangle}_{t-1}).$ In practice, we often recommend this choice.
  • A typical choice we recommend is $M=[5,\cdots,20]$. The performance with varying $M$ values will be discussed in Section 6.
  • Line 3 can be computed in parallel at the expense of additional memory.
  • When the normalizing constant can be calculated relatively easily (e.g., in discrete diffusion with small $K,L$), we can directly sample from $\{p^{\star,\alpha}_{t}\}_{t=T}^{0}$.
  • We can also use a proposal distribution different from $p^{\mathrm{pre}}_{t-1}$. For the extension, refer to Section C.
The remaining question is how to obtain the soft value function, which we address in the next section.
4.3 Learning Soft Value Functions
Next, we describe how to learn soft value functions $v_{t}(x)$ in practice. We propose two main approaches: the Monte Carlo regression approach and the posterior mean approximation approach.
Monte Carlo regression.
Here, we use the following approximation $v^{\prime}_{t}$ as $v_{t}$ where
$v^{\prime}_{t}(\cdot):=\mathbb{E}_{x_{0}\sim p^{\mathrm{pre}}(x_{0}|x_{t})}[r(x_{0})|(x_{t})=\cdot].$
This is based on
$v_{t}(x_{t}):=\alpha\log\mathbb{E}_{x_{0}\sim p^{\mathrm{pre}}(x_{0}|x_{t})}[\exp(r(x_{0})/\alpha)|x_{t}]\approx\log(\exp(\mathbb{E}_{x_{0}\sim p^{\mathrm{pre}}(x_{0}|x_{t})}[r(x_{0})|x_{t}]))=v^{\prime}_{t}(x_{t}).$ (1)
By regressing $r(x_{0})$ onto $x_{t}$, we can learn $v^{\prime}_{t}$ as in Algorithm 2. Combining this with Algorithm 1, we refer to the entire optimization approach as SVDD-MC.
Algorithm 2 Value Function Estimation Using Monte Carlo Regression
1:  Require: Pre-trained diffusion models, reward $r:\mathcal{X}\to\mathbb{R}$, function class $\Phi:\mathcal{X}\times[0,T]\to\mathbb{R}$.
2:  Collect datasets $\{x^{(s)}_{T},\cdots,x^{(s)}_{0}\}_{s=1}^{S}$ by rolling-out $\{p^{\mathrm{pre}}_{t}\}_{t=T}^{0}$ from $t=T$ to $t=0$.
3:  $\hat{v}^{\prime}=\mathop{\mathrm{argmin}}_{f\in\Phi}\sum_{t=0}^{T}\sum_{s=1}^{S}\{r(x^{(s)}_{0})-f(x^{(s)}_{t},t)\}^{2}.$
4:  Output: $\hat{v}^{\prime}$
Note that technically, without approximation introduced in (1), we can estimate $v_{t}$ by regressing $\exp(r(x_{0})/\alpha)$ onto $x_{t}$ based on the original definition. This approach may work in many cases. However, when $\alpha$ is very small, the scaling of $\exp(r(\cdot)/\alpha)$ tends to be excessively large. Due to this concern, we generally recommend using using Algorithm 2.
Remark 2 (Another way of constructing datasets).
We have used the backward process in pre-trained diffusion models to collect training datasets in Algorithm 2. Given that the backward and forward processes are designed to match marginally during the training of diffusion models as discussed in Section 3.1, we can also use the forward process in pre-trained diffusion models to construct datasets.
Remark 3 (Another way of learning value functions).
Technically, another method for learning value functions is available such as soft-Q-learning (Section D), by leveraging soft-Bellman equations in diffusion models Uehara et al. (2024, Section 3) (as known in standard RL contexts (Levine, 2018; Haarnoja et al., 2017; Geist et al., 2019)). However, since we find Monte Carlo approaches to be more stable, we recommend them over soft-Q-learning.
Posterior mean approximation.
Here, recalling we use $\hat{x}_{0}(x_{t})$ (approximation of $\mathbb{E}_{x_{0}\sim p^{\mathrm{pre}}(x_{t})}[x_{0}|x_{t}]$) when training pre-trained diffusion models in Section 3.1, we perform the following approximation:
$v_{t}(x):=\alpha\log\mathbb{E}_{x_{0}\sim p^{\mathrm{pre}}(x_{0}|x_{t})}[\exp(r(x_{0})/\alpha)|x_{t}]\approx\alpha\log(\exp(r(\hat{x}_{0}(x_{t}))/\alpha)=r(\hat{x}_{0}(x_{t})).$
Then, we can use $r(\hat{x}_{0}(x_{t}))$ as the estimated value function.
The advantage of this approach is that no additional training is required as long as we have $r$. When combined with Algorithm 1, we refer to the entire approach as SVDD-PM.
Algorithm 3 Value Function Estimation using Posterior Mean Approximation
1:  Require: Pre-trained diffusion models, reward $r:\mathcal{X}\to\mathbb{R}$
2:  Set $\hat{v}^{\diamond}(\cdot,t):=r(\hat{x}_{0}(x_{t}=\cdot),t)$
3:  Output: $\hat{v}^{\diamond}$
Remark 4 (Relation with DPS).
In the context of classifier guidance, similar approximations have been employed (e.g., DPS in Chung et al. (2022)). However, the final inference-time algorithms differ significantly, as these methods compute gradients at the end.
5 Advantages, Limitations, and Extensions of SVDD
So far, we have detailed our algorithm, SVDD. In this section, we discuss its advantages, limitations, and extensions.
5.1 Advantages
No fine-tuning (or no training in SVDD-PM).
Unlike classifier-free guidance or RL-based fine-tuning, SVDD  does not require any fine-tuning of the generative models. In particular, when using SVDD-PM, no additional training is needed as long as we have $r$.
No need for constructing differentiable models.
Unlike classifier guidance, SVDD  does not require differentiable proxy models, as there is no need for derivative computations. For example, if $r$ is non-differentiable feedback (e.g., physically-based simulations, QED, SA in molecule generation), our method SVDD-PM can directly utilize such feedback without constructing differentiable proxy models. In cases where non-differentiable feedback is costly to obtain, proxy reward models may still be required, but they do not need to be differentiable; thus, non-differentiable features or non-differentiable models based on scientific knowledge (e.g., molecule fingerprints, GNNs) can be leveraged. Similarly, when using SVDD-MC, while a value function model is required, it does not need to be differentiable, unlike classifier guidance.
Additionally, compared to approaches that involve derivatives (like classifier guidance or DPS), our algorithm may be more memory efficient at inference time, particularly when $M$ is moderate, and can be directly applied to discrete diffusion models mentioned in Example 2.
Proximity to pre-trained models (robust to reward over-optimization).
Since samples are consistently generated from pre-trained diffusion policies at each step, they are ensured to remain within the natural space defined by pre-trained diffusion models.
This is especially advantageous when rewards to be optimized by SVDD are learned from offline data. In such cases, learned reward functions may be inaccurate for out-of-distribution samples (Uehara et al., 2024). Consequently, conventional fine-tuning methods often suffer from over-optimization (reward hacking) by exploiting these out-of-distribution regions (Fan et al., 2023; Clark et al., 2023). Given that non-natural design spaces typically encompass a significant portion of out-of-distribution regions of the offline data, maintaining proximity to pre-trained models acts as a regularization mechanism against these regions.
5.2 Potential Limitations
Memory and computational complexity in inference time.
Our approach requires more computational resources (if not parallelized) or memory (if parallelized), specifically (roughly) $M$ times more than standard inference methods. Considering this aspect, we conduct a fair comparison between SVDD and best-of-N in Section 6.
Proximity to pre-trained models.
The proximity to pre-trained models might be a disadvantage if significant changes to the pre-trained models are desired. We acknowledge that RL-based fine-tuning could be more effective for this purpose than our algorithm.
5.3 Extensions
Using a likelihood/classifier as a reward.
While we primarily consider scenarios where reward models are regression models, by adopting a similar strategy in (Zhao et al., 2024), they can be readily replaced with classifiers or likelihood functions in the context of solving inverse problems (Chung et al., 2022; Bansal et al., 2023).
Application to fine-tuning.
Our SVDD, can also be naturally extended to fine-tuning by generating samples with SVDD, and then using these samples for supervised fine-tuning.
5.4 Comparison between SVDD and Standard SMC-Methods
The algorithm most closely related to SVDD, are several SMC-Methods for guidance (Wu et al., 2024; Trippe et al., 2022; Dou and Song, 2024; Phillips et al., 2024; Cardoso et al., 2023), which utilizes “standard” sequential Monte Carlo (Del Moral and Doucet, 2014; Kitagawa, 1993) in diffusion models. Although originally designed for conditioning, it can technically be applied to reward maximization. However, they require resampling across the “entire” batch, complicating parallelization. Additionally, when batch sizes are small, as is often the case in recent large diffusion models, its performance may be suboptimal. In contrast, our method is easily parallelizable and can achieve strong performance even with a batch size of $1$, as resampling occurs for each individual sample. For further details, see Section A.
6 Experiments
We conduct experiments to assess the performance of our algorithm relative to baselines and its sensitivity to various hyperparameters. We start by outlining the experimental setup, including baselines and models, and then present the results.
6.1 Settings
Methods to compare.
We compare the following methods.
  • Pre-trained models: We generate samples using pre-trained models.
  • Best-of-N (Nakano et al., 2021): We generate samples from pre-trained models and select the top $1/N$ samples. This selection is made to ensure that the computational time during inference is approximately equivalent to that of our proposal.
  • DPS (Chung et al., 2022): DPS is a widely used enhancement of classifier guidance. Although the original work was not designed for discrete diffusion, we employ specific heuristics to adapt it for this purpose (Section E).
  • SVDD (Ours): We implement SVDD-MC and SVDD-PM. We generally set $M=20$ for images and $M=10$ for other domains, additionally we set $\alpha=0$.
Datasets and reward models.
We provide details on the pre-trained diffusion models and downstream reward functions used. For further information, refer to Section E.
  • Images: We use Stable Diffusion v1.5 as the pre-trained diffusion model ($T=50$). For downstream reward functions, we use compressibility and aesthetic scores (LAION Aesthetic Predictor V2 in Schuhmann (2022)), as employed by Black et al. (2023); Fan et al. (2023). Note that compressibility is non-differentiable reward feedback.
  • Molecules: We use GDSS (Jo et al., 2022), trained on ZINC-250k (Irwin and Shoichet, 2005), as the pre-trained diffusion model ($T=1000$). For downstream reward functions, we use QED and SA calculated by RDKit, which are non-differentiable feedback. Here, we renormalize SA to $(10-\mathrm{SA})/9$ so that a higher value indicates better performance.
  • DNAs (Enhancers): We use the discrete diffusion model (Sahoo et al., 2024), trained on datasets from Gosai et al. (2023), as our pre-trained diffusion model ($T=128$). For the downstream reward function, we use an Enformer model (Avsec et al., 2021) to predict activity in the HepG2 cell line.
  • RNAs (5’UTRs): We use the discrete diffusion model (Sahoo et al., 2024) as the pre-trained diffusion model ($T=128$). For downstream reward functions, we employ a reward model that predicts the mean ribosomal load (MRL) measured by polysome profiling (trained on datasets from Sample et al. (2019)) and stability (trained on datasets from (Agarwal and Kelley, 2022)).
6.2 Results
(a) Images: compressibility
(b) Images: aesthetic score
(c) Molecules: QED
(d) Molecules: SA
(e) Enhancers
(f) 5’UTRs: MRL
(g) 5’UTRs: stability
Figure 2: We show the histogram of generated samples in terms of reward functions.
(a) Images: compressibility
(b) Images: aesthetic scores
(c) Molecules: QED scores
(d) Molecules: SA scores (Normalized as $(10-SA)/9$ )
Figure 3: We show generated samples from our proposal. For more samples, refer to Section E.2.
We compare the baselines with our two proposed methods. The performance is shown in Figure 2, and the generated samples are presented in Figure 3.
Overall, our proposal outperforms the baseline methods (Best-of-N and DPS), as evidenced by higher rewards for samples generated in large quantities. In Section E, we compare the top $10^{\text{th}}$ and $20^{\text{th}}$ quantiles from each algorithm and confirm that SVDD always outperforms the baselines. This indicates that our algorithm can generate high-reward samples that Best-of-N and DPS cannot.
  • Compared to Best-of-N, although the rewards for generating samples in smaller quantities could be lower with our algorithm, this is expected because our algorithm generates samples with high likelihood $p^{\mathrm{pre}}(x|c)$ in pre-trained diffusion models, but with possibly lower rewards.
  • Compared to DPS, our algorithm consistently achieves better performance. Notably, in molecular generation, DPS is highly ineffective due to the non-differentiable nature of the original feedback.
The relative performance of our two proposed methods (SVDD-MC or SVDD-PM) appears to be domain-dependent. Generally, SVDD-PM may be more robust since it does not require additional learning (i.e., it directly utilizes reward feedback). The performance of SVDD-MC depends on the success of value function learning, which is discussed in Section E.
Figure 4: Performance of our algorithm SVDD as $M$ varies, for image generation while optimizing the aesthetic score.
Ablation studies in terms of $M$.
We plot the performance as $M$ varies. The performance gradually reaches a plateau as $M$ increases. This tendency is seen in all other domains.
7 Conclusion
We propose a novel inference-time algorithm, SVDD, for optimizing downstream reward functions in pre-trained diffusion models that eliminate the need to construct differentiable proxy models. In future work, we plan to conduct experiments in other domains, such as protein sequence optimization (Gruver et al., 2023; Alamdari et al., 2023; Watson et al., 2023) or controllable 3D molecule generation (Xu et al., 2023).
Acknowledgement
We appreciate the feedback from Christian A. Naesseth. He pointed out the connection between our work and the twisted diffusion sampler, and noted our algorithm is an instantiation of nested-IS SMC in our context (Naesseth et al., 2019, 2015).
Appendix A Difference Between SVDD and “Standard” SMC-Methods
Algorithm 4 Guidance with “Standard” SMC (for reward maximization)
1:  Require: Estimated value functions $\{\hat{v}_{t}(x)\}_{t=T}^{0}$, pre-trained diffusion models $\{p^{\mathrm{pre}}_{t}\}_{t=T}^{0}$, hyperparameter $\alpha\in\mathbb{R}$, Batch size $N$
2:  for $t\in[T+1,\cdots,0]$ do
3:     IS step:
4:
$i\in[1,\cdots,N];x^{[i]}_{t-1}\sim p^{\mathrm{pre}}_{t-1}(\cdot|x^{[i]}_{t}),w^{[i]}_{t-1}:=\frac{\exp(\hat{v}_{t-1}(x^{[i]}_{t-1})/\alpha)}{\exp(\hat{v}_{t}(x^{[i]}_{t})/\alpha)}$
5:     Selection step: select new indices with replacement
6:     $\{x^{[i]}_{t-1}\}_{i=1}^{N}\leftarrow\{x^{\zeta^{[i]}_{t-1}}_{t-1}\}_{i=1}^{N},\quad\{\zeta^{[i]}_{t-1}\}_{i=1}^{N}\sim\mathrm{Cat}\left(\left\{\frac{w^{[i]}_{t-1}}{\sum_{j=1}^{N}w^{[j]}_{t-1}}\right\}_{i=1}^{N}\right)$
7:  end for
8:  Output: $x_{0}$
In this section, we compare our algorithm with the SMC-based methods (Wu et al., 2024; Trippe et al., 2022; Dou and Song, 2024; Phillips et al., 2024; Cardoso et al., 2023) for guidance. While they originally aim to solve a conditioning problem, which is different from reward maximization, they algorithms can be converted to reward maximization. We then elaborate on the differences between SVDD, and them in the context of reward maximization. Notably, our SVDD is an instantiation of nested-IS SMC (Naesseth et al., 2019, Algorithm 5) in the literature on computational statistics, whereas these SMC-based Methods rely on standard sequential Monte Carlo.
A.1 SMC-Based Methods
The complete algorithm of TDS in our setting is summarized in Algorithm 4. Since our notation and their notations are slightly different, we first provide a brief overview. It consists of two steps. Since our algorithm is iterative, at time point $t$, consider we have $N$ samples (particles) $\{x^{[i]}_{t}\}_{i=1}^{N}$.
IS step (line 3).
We generate a set of samples $\{x^{[i]}_{t-1}\}_{i=1}^{N}$ following a policy from a pre-trained model $p^{\mathrm{pre}}_{t-1}(\cdot|\cdot)$. In other words,
$\forall i\in[1,\cdots,N];x^{[i]}_{t-1}\sim p^{\mathrm{pre}}_{t-1}(\cdot|x^{[i]}_{t}).$
Now, we denote the importance weight for the next particle $x_{t-1}$ given the current particle $x_{t}$ as $w(x_{t-1},x_{t})$, expressed as
[…]
and define
$\forall i\in[1,\cdots,N];\quad w^{[i]}_{t-1}:=w(x^{[i]}_{t-1},x^{[i]}_{t}).$
Note here we have used the soft Bellman equation:
[…]
Hence, by denoting the target marginal distribution at $t-1$, we have the following approximation:
$p^{\mathrm{tar}}_{t-1}\underbrace{\approx}_{\text{IS}}\sum_{i=1}^{N}\frac{w^{[i]}_{t-1}}{\sum_{j=1}^{N}w^{[j]}_{t-1}}\delta_{x^{[i]}_{t-1}}.$
Selection step (line 5).
Finally, we consider a resampling step. The resampling indices are determined by the following:
$\{\zeta^{[i]}_{t-1}\}_{i=1}^{N}\sim\mathrm{Cat}\left(\left\{\frac{w^{[i]}_{t-1}}{\sum_{j=1}^{N}w^{[j]}_{t-1}}\right\}_{i=1}^{N}\right).$
To summarize, we conduct
$p^{\mathrm{tar}}_{t-1}\underbrace{\approx}_{\text{IS}}\sum_{i=1}^{N}\frac{w^{[i]}_{t-1}}{\sum_{j=1}^{N}w^{[j]}_{t-1}}\delta_{x^{[i]}_{t-1}}\underbrace{\approx}_{\text{Resampling}}\frac{1}{N}\sum_{i=1}^{N}\delta_{x^{\zeta_{t-1}^{[i]}}_{t-1}}.$
Finally, we give several important remarks.
  • All of existing works Wu et al. (2024); Cardoso et al. (2023); Phillips et al. (2024); Dou and Song (2024) actually consider a scenario where the reward $r$ is a classifier. In Algorithm 4, we tailor the algorithm for reward maximization. Vice verisa, as we mentioned in Section 5.3, our SVDD can also operate effectively when $r$ is a classifier.
  • In Wu et al. (2024); Cardoso et al. (2023); Phillips et al. (2024), the proposal distribution is not limited to the pre-trained model. Likewise, in our SVDD, we can select an arbitrary proposal distribution, as discussed in Section C.
  • In the context of autoregressive (language) models, Zhao et al. (2024); Lew et al. (2023) proposed a similar algorithm.
A.2 Comparison of SVDD with “Standard” SMC-Based Methods (SSM) for Reward Maximization
We now compare our SVDD with “standard” SMC-Based Methods (SSM). Here, we write a batch size of SVDD in $G$. Importantly, we note that our implementation is analogous to nested-IS SMC in the literature in computational statistics; hence, many differences between nested-IS SMC (Naesseth et al., 2019, 2015) and pure SMC in computational statistics are translated here.
We first reconsider the fundamental assumptions of each algorithm. SVDD’s performance, in terms of rewards, depends on the size of $M$ but is independent of the batch size $G$. In contrast, the performance of SSM depends on the batch size $N$. With this in mind, we compare the advantages of SVDD over SSM from various perspectives.
Ease of parallelization in SVDD.
SVDD is significantly easier to parallelize across multiple nodes. In contrast, SSM requires interaction between nodes for parallelization. This advantage is also well-documented in the context of nested-IS SMC versus standard SMC (Naesseth et al., 2019).
High performance of SVDD under memory constraints.
Now, consider a scenario where the batch size is small, which often occurs due to memory constraints in large pre-trained diffusion models. In this case, while SSM may exhibit suboptimal performance, SVDD-PM can still achieve high performance by choosing a sufficiently large $M$ (assuming that models of the value functions are not significantly larger than those in pre-trained diffusion models).
In SSM, the “ratio” is approximated.
In SVDD, we approximate each $\exp(v_{t-1}(x_{t-1})/\alpha)$ as a weight. However, in standard SMC, the ratio is approximated as a weight:
$\frac{\exp(v_{t-1}(x_{t-1})/\alpha)}{\exp(v_{t}(x_{t})/\alpha)}.$
The key difference is that in SSM, both the numerator and the denominator are approximated, which could lead to greater error propagation.
Appendix B Comparising against DG (Nisonoff et al., 2024) and DiGress (Vignac et al., 2022) in Discrete Diffusion Models
Our method is closely related to guidance methods used in DG (Nisonoff et al., 2024) and DiGress (Vignac et al., 2022). However, we emphasize that our approach is more general, as it operates on any domain, including continuous spaces including Riemannian spaces and discrete spaces, in a unified manner, making a direct comparison not feasible. With this in mind, we provide a comparison focusing on cases where all domains are discrete.
Comparison with Nisonoff et al. (2024).
In this continuous framework, following the notation in (Lou et al., 2023), they propose the use of the following rate matrix:
$Q^{\star}_{x,y}(t)=Q^{\mathrm{pre}}_{x,y}(t)\frac{\exp(v_{t}(y)/\alpha)}{\exp(v_{t}(x)/\alpha)}.$
where $Q^{\mathrm{pre}}_{x,y}(t)$ is a rate matrix in the pre-trained model. This suggests that, with standard discretization, the optimal policy at each time step is:
$p(x_{t+\delta t}=y|x_{t}=x)=\mathrm{I}(x\neq y)+Q^{\star}_{x,y}(t)(\delta t)$ (2)
where $\delta t$ is step size. Asymptotically, this is equivalent to sampling from the optimal policy in Theorem 1, as we will show in Remark 5. However, as (Nisonoff et al., 2024) note, sampling from the optimal policy requires $O(KL)$ computation, where $K$ is the vocabulary size and $L$ is the sequence length, which is computationally expensive for large $K$ and $L$. To address this issue, they propose using a Taylor approximation by computing the gradient once. In contrast, we avoid this computational overhead in a different manner, i.e., through importance sampling and resampling. A detailed empirical comparison between our method and theirs is reserved for future work.
Remark 5 (Asymptotic equivalence between formula in Nisonoff et al. (2024) and Theorem 1 ).
The informal reasoning is as follows. Recall that the pre-trained policy can be written as
$p(x_{t+\delta t}=y|x_{t}=x)=\mathrm{I}(x\neq y)+Q^{\mathrm{pre}}_{x,y}(t)(\delta t).$
Then, Theorem 1 states that the optimal policy is
$\frac{\{\mathrm{I}(x\neq y)+Q^{\mathrm{pre}}_{x,y}(t)(\delta t)\}\exp(v_{t}(y)/\alpha)}{\sum_{z}\{\mathrm{I}(x\neq z)+Q^{\mathrm{pre}}_{x,z}(t)(\delta t)\}\exp(v_{t}(z)/\alpha)}.$
Now, we have
$\frac{\{\mathrm{I}(x\neq y)+Q^{\mathrm{pre}}_{x,y}(t)(\delta t)\}\exp(v_{t}(y)/\alpha)}{\sum_{z}\{I(x\neq z)+Q^{\mathrm{pre}}_{x,z}(t)(\delta t)\}\exp(v_{t}(z)/\alpha)}$
$=\frac{\mathrm{I}(x\neq y)+Q^{\mathrm{pre}}_{x,y}(t)(\delta t)\}\exp(v_{t}(y)/\alpha)}{\exp(v_{t}(x)/\alpha)}\times\left\{1+O(\delta t)\right\}$
$\approx\frac{\{\mathrm{I}(x\neq y)+Q^{\mathrm{pre}}_{x,y}(t)(\delta t)\}\exp(v_{t}(y)/\alpha)}{\exp(v_{t}(x)/\alpha)}=\mathrm{I}(x\neq y)+\frac{Q^{\mathrm{pre}}_{x,y}(t)(\delta t)\exp(v_{t}(y)/\alpha)}{\exp(v_{t}(x)/\alpha)}.$
Thus, this recovers the formula (2).
Comparison with DiGress in Vignac et al. (2022).
Vignac et al. (2022) proposed a diffusion model for graphs where each sampling and denoising step operates directly on the discrete structure, avoiding continuous relaxation. They discuss how to implement guidance by treating rewards as a classifier. To bypass the exponential computational cost of sampling from the optimal policy ($p^{(\alpha)}$ in Theorem 1), they employ a Taylor expansion. While it requires the calculation of gradients for value functions, then it mitigates the exponential blow-up in computational time. In contrast, we avoid this computational blow-up through importance sampling (IS) and resampling. A detailed empirical comparison between our method and theirs is left for future work.
Remark 6.
Note that in our molecule generation experiment in Section 6, we use GDSS, which operates in continuous space and differs from DiGress.
Appendix C Extension with Arbitrary Proposal Distribution
Here, we describe the algorithm where the proposal distribution is not necessarily derived from the policy of the pre-trained model, as summarized in Algorithm 5. Essentially, we only adjust the importance weight. In practice, we can use the gradient of a differentiable proxy model, such as DPS, as the proposal distribution $q_{t-1}$. Even if the differentiable proxy (value function) models are not highly accurate, our method will still perform effectively since other value function models $\hat{v}_{t-1}$ can be non-differentiable.
Algorithm 5 SVDD (Soft Value-Based Decoding in Diffusion Models)
1:  Require: Estimated soft value function $\{\hat{v}_{t}\}_{t=T}^{0}$ (refer to Algorithm 2 or Algorithm 3), pre-trained diffusion models $\{p^{\mathrm{pre}}_{t}\}_{t=T}^{0}$, hyperparameter $\alpha\in\mathbb{R}$, proposal distribution $\{q_{t}\}_{t=T}^{0}$
2:  for $t\in[T+1,\cdots,1]$ do
3:     Get $M$ samples from pre-trained polices $\{x^{\langle m\rangle}_{t-1}\}_{m=1}^{M}\sim q_{t-1}(\cdot|x_{t})$, and for each $m$, and calculate
$w^{\langle m\rangle}_{t-1}:=\exp(\hat{v}_{t-1}(x^{\langle m\rangle}_{t-1})/\alpha)\times\frac{p^{\mathrm{pre}}_{t-1}(x^{\langle m\rangle}_{t-1}|x_{t})}{q_{t-1}(x^{\langle m\rangle}_{t-1}|x_{t})}.$
4:     $x_{t-1}\leftarrow x^{\langle\zeta_{t-1}\rangle}_{t-1}$ after selecting an index: $\zeta_{t-1}\sim\mathrm{Cat}\left(\left\{\frac{w^{\langle m\rangle}_{t-1}}{\sum_{j=1}^{M}w^{\langle j\rangle}_{t-1}}\right\}_{m=1}^{M}\right),\,$
5:  end for
6:  Output: $x_{0}$
Appendix D Soft Q-learning
In this section, we explain soft value iteration to estimate soft value functions, which serves as an alternative to Monte Carlo regression.
Soft Bellman equation.
Here, we use the soft Bellman equation:
[…]
as proved in Section 4.1 in (Uehara et al., 2024). In other words,
$v_{t}(x_{t})=\alpha\log\{\mathbb{E}_{x_{t-1}\sim p^{\mathrm{pre}}(\cdot|x_{t})}\left[\exp(v_{t-1}(x_{t-1})/\alpha)|x_{t}\right]\}.$
Algorithm.
Based on the above, we can estimate soft value functions recursively by regressing $v_{t-1}(x_{t-1})$ onto $x_{t}$. This approach is often referred to as soft Q-learning in the reinforcement learning literature (Haarnoja et al., 2017; Levine, 2018).
Algorithm 6 Value Function Estimation Using Soft Q-learning
1:  Require: Pre-trained diffusion models $\{p^{\mathrm{pre}}_{t}\}^{0}_{t=T}$, value function model $v(x;\theta)$
2:  Collect datasets $\{x^{(s)}_{T},\cdots,x^{(s)}_{0}\}_{s=1}^{S}$ by rolling-out $\{p^{\mathrm{pre}}_{t}\}_{t=T}^{0}$ from $t=T$ to $t=0$.
3:  for $j\in[0,\cdots,J]$ do
4:     Update $\theta$ by running regression:
$\theta^{\prime}_{j}\leftarrow\mathop{\mathrm{argmin}}_{\theta}\sum_{t=0}^{T}\sum_{s=1}^{S}\left\{v(x^{(s)}_{t};\theta)-v(x^{(s)}_{t-1};\theta^{\prime}_{j-1})\right\}^{2}.$
5:  end for
6:  Output: $v(x;\theta^{\prime}_{J})$
In our context, due to the concern of scaling of $\alpha$, as we have done in Algorithm 2, we had better use
$v_{t}(x_{t})=\mathbb{E}_{x_{t-1}\sim p^{\mathrm{pre}}(\cdot|x_{t})}\left[v_{t-1}(x_{t-1})|x_{t}\right].$
With the above recursive equation, we can estimate soft value functions as in Algorithm 6.
Appendix E Additional Experimental Details
We further add additional experimental details.
E.1 Additional Setups for Experiments
Images.
  • DPS: We require differentiable models that map images to compressibility. For this task, we have used a standard CNN.
Molecules.
  • DPS: Following the implementation in Lee et al. (2023), we use the same GNN model as the reward model. Note that this model cannot compute derivatives with respect to adjacency matrices.
  • SVDD-MC: We use a Graph Isomorphism Network (GIN) model (Xu et al., 2018) as a value function model.
Enhancers.
  • DPS: Although DPS was originally proposed in the continuous space, we have adapted it for our use by incorporating the gradient of the value function model at each step and representing each sequence as a one-hot encoding vector.
  • SVDD-MC: We have used the Enformer model as the function model.
5’UTRs.
We have used ConvGRU as the reward model (Dey and Salem, 2017).
  • DPS: Although DPS was originally proposed in the continuous space, we have adapted it for our use by incorporating the gradient of the value function model at each step and representing each sequence as a one-hot encoding vector.
  • SVDD-MC: We employed ConvGRU as the value function model.
E.2 Additional Results
We evaluate the performance of the generated samples using histograms in Section 6. Here, we present the top 10 and 20 quantiles of the generated samples in Table 2.
Table 2: Top 10 and 20 quantiles of the generated samples for each algorithm. Higher is better.
Domain Quantile Pre-Train Best-N DPS SVDD-MC SVDD-PM
Images compress 20% -86.2 -63.2 -67.1 - -43.7
10% -78.6 -57.3 -61.2 - -38.8
Images aesthetic 20% 5.875 6.246 5.868 - 6.356
10% 5.984 6.343 5.997 - 6.472
Molecules QED 20% 0.771 0.881 0.802 0.912 0.916
10% 0.812 0.902 0.843 0.925 0.928
Molecules SA 20% 0.750 0.916 0.802 1.0 1.0
10% 0.803 0.941 0.849 1.0 1.0
Enhancers 20% 0.74 3.00 2.68 5.53 6.44
10% 1.41 3.52 3.85 5.75 7.02
5’UTR MRL 20% 0.78 0.97 0.90 1.09 1.33
10% 0.86 1.021 0.93 1.12 1.38
5’UTR Stability 20% -0.63 -0.59 -0.62 -0.52 -0.56
10% -0.61 -0.58 -0.60 -0.51 -0.55
Performance of value function training.
We report the performance of value function learning using Monte Carlo regression as follows. We plot the Pearson correlation on the test dataset.
(a) Images: compressibility
(b) Images: aesthetic score
(c) Molecules: QED
Figure 5: Training curve
More generated samples.
We have provided additional generated samples in Figure 6, Figure 7, Figure 8, Figure 9, and Figure 10.
Figure 6: Additional generated samples (Domain: images, Reward: Compressibility)
Figure 7: Additional generated samples (Domain: Images, Reward: Aesthetic score)
Figure 8: Additional generated samples (Domain: Molecules, Reward: QED score)
Figure 9: Additional generated samples (Domain: Molecules, Reward: SA score, normalized as $(10-SA)/9$)
Figure 10: Additional generated samples from SVDD-MC (Domain: Molecules, Reward: SA score = 1.0 (normalized as $(10-SA)/9$))